OpenStudy (anonymous):

I have a story problem where I have to come up with 1/3 of 1.8x10 to the 24th power. i am assuming that I need to divide each side by the third but I need to simplify in scientific notation. Do I just divide each side and leave the 24th power or do I divide the power as well by 1/3?

OpenStudy (adamaero):

Use "equation" tool. It's hard to understand what equation

OpenStudy (anonymous):

\[Sorry didn't know about that. Thank You. I need 1\3\left( 1.8\times 10^{24}\right)\]

OpenStudy (anonymous):

sorry again trying to get the hang of the tool.

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }of \left( 1.8\times10^{24} \right)\]